OpenStudy (anonymous):

In a box of 75 markers, 36 matkers are either red or black, and 15 are blue. What is the probability that a randomly selected marker is red or black or blue. I have this 4 options: A. 32 B. 68 C. 51 D. 24

OpenStudy (anonymous):

The markers are either red, black or blue aren't they? So when you pick one you would definately get one of the three right? That would be a probability of 1 or 100%... Are you sure you've asked the question correctly? Probabilities are always within a range of 0 and 1.

OpenStudy (anonymous):

Oh wait, there are 75 total and only 51 that are red/black/blue but the probabilities still don't make sense.

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

That's exactly how the exercise is given to me

OpenStudy (anonymous):

Well it's probably meant as a percentage now that I look at it. So there are 75 markers total and 51 are red/blue/black. 51/75 = 68%

OpenStudy (anonymous):

They give me 4 ootion 32 68 51 24

OpenStudy (anonymous):

Oh okay thankssssssss :D

OpenStudy (anonymous):

This is true because red/blue/black make up 68 percent of all of the markers. So when you randomly choose one your odds are 68% that it is one of those.

OpenStudy (callisto):

possible outcome = 36+15 = 51 total outcome = 75 probability = possible outcome / total outcome =... Probability cannot be greater than 1...

OpenStudy (callisto):

of course, in this case, you have to pick up the most sensible choice from these non sensible choices... If I were you , i would choose 51, then complain to the teacher :P

OpenStudy (anonymous):

:)